OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. my equation is F(x) = 1 (x-10)^2 + 8 in vertex form and and y = x^2 - 20x + 108. Any ideas how to do this would be super helpful and I'll give you a medal!

OpenStudy (ranga):

The average rate of a function f(x) in the interval x = a to x = b is defined as: \[\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

HOW DO YOU FIND A AND B

OpenStudy (ranga):

a and b are just any two points. Here, a = 4 and b = 7. \[\frac{f(7)-f(4)}{7-4} = \frac{f(7)-f(4)}{3}\]Here: f(x) = x^2 - 20x + 108 f(7) = ? f(4) = ? Plug the numbers into the average rate formula above.

OpenStudy (anonymous):

OOOOO Ok thank you so much

OpenStudy (ranga):

You are welcome.